. A chemical plant stores spare parts for maintenance in a large warehouse. Throughout the working day, maintenance personnel go to the warehouse to pick up supplies needed for their jobs. The warehouse receives a request for supplies, on average, every three minutes. The average request requires 2.75 minutes to fill a request. Maintenance employees are paid $21.50 per hour and warehouse employees are paid $16 per hour. The warehouse operates 8 hours per day.

a) Based on the number of maintenance employees in the system, an 8 hour work day, and the given arrival and service rates. What is the system cost per day (to the nearest $) if there is only 1 warehouse employees working?

b) Based on the number of maintenance employees in the system, an 8 hour work day, and the given arrival and service rates. What is the system cost per day (to the nearest $) if there are 2 warehouse employees working?

c) Based on the number of maintenance employees in the system, an 8 hour work day, and the given arrival and service rates. What is the system cost per day (to the nearest $) if there are 3 warehouse employees working?

d) What is the optimal number of warehouse employees to staff the warehouse?
Answer #1

a) The number of maintenance employees is not specified in the problem, so consider that it is an infinite number. So, it is a M/M/1 system with following parameters,

Arrival rate, \lambda = 1/interarrival time = 1/3 per minute = (1/3)*60 = 20 per hour

Service rate, \mu = 1/service time = (1/2.75)*60 = 21.82 per hour

Number of maintenance employees in the system, L = \lambda/(\mu-\lambda) = 20/(21.82-20) = 11

System cost per hour = Cw*L + s*Cs

= 21.5*11+1*16

= $ 252.5

System cost per day = 252.5*8

= $ 2,020

b) With more than 1 warehouse employees (servers) working, it is a multi-channel M/M/s system , where

Number of servers, s = 2

Operating characteristics are computed using spreadsheet as below

D11 Q fx =(VLOOKUP(D6+1,$GS2SH$26,2,0)-VLOOKUPfD6,SG$2:SH$26,2,0))D9D10/(1-D10)^2+D4/D5 1 A Mutliple-Server Waiting Line Sy

EXCEL FORMULAS:

Cell Formula H3 (SDS4/SDS5)AG2/FACT(G2)+H2 D9 Copy to H3:H26 1/((VLOOKUP(D6+1,SGS2:SHS26,2,0)-VLOOKUP(D6,SGS2:SH$26,2,0)/(1 D

L = 1.1604

Total system cost per hour = 1.1604*21.5+2*16 = 56.95

Total system cost per day = 56.95*8 = $ 455.6

b) With more than 1 warehouse employees (servers) working, it is a multi-channel M/M/s system , where

Number of servers, s = 3

Operating characteristics are computed using spreadsheet as below:

n Summation 1 A Mutliple-Server Waiting Line System 2 2 Input Arrival rate- per hour 2 2 2 Service rate21.818182 per hour No.

L = 0.9489

Total system cost per hour = 0.9489*21.5+3*16 = 68.40

Total system cost per day = 68.40*8 = $ 547.2

d) We see that total system cost per day of 2 warehouse employees working is the lowest.

Therefore, optimal number of warehouse employees to staff the warehouse = 2